OpenStudy (anonymous):

Solve the following system of equations: −2x + y = 1 −4x + y = −1 (3, 1) (−1, 3) (−1, −3) (1, 3)

OpenStudy (anonymous):

You can use a method known as elimination, in which you eliminate one of the variables in order to solve for the other one. In this case, it seems easiest to eliminate y and solve for x. Eliminate y using subtraction: −2x + y = 1 −4x + y = −1 - ________________ 2x = 2 x=1 Plug in x to solve for y: -2(1) + y = 1 y = 3

OpenStudy (anonymous):

(1,3)